OpenStudy (anonymous):

What is the area of a rectangle with a length of 12 and a perimeter of 30?

OpenStudy (anonymous):

p=2(l+w) w=(p/2)-l A=l*w

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

oh ok so it's 2(12+30)

OpenStudy (anonymous):

No Perimeter = 2(length + breadth)

OpenStudy (anonymous):

Find breadth and Area = length x breadth

OpenStudy (anonymous):

whats breadth

OpenStudy (anonymous):

w=(30/2)-12 w=3 A=12*3 A=36

OpenStudy (anonymous):



OpenStudy (anonymous):

It is also called width

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome